OpenStudy (anonymous):

This is a differential equation of 2nd order, I need to find a function that if substituted on both sides will be equal

OpenStudy (anonymous):

\[\frac{ d ^{2} x }{ dt ^{2} }= -\omega ^{2}x\]

OpenStudy (anonymous):

You have constant coefficients in the homogeneous equation, \[\frac{d^2x}{dt^2}+\omega^2x=0\] which has the characteristic equation, \[r^2+\omega^2=0\] with roots \(r=\pm\sqrt{-\omega}=\pm i\omega\). The general solution would then be \[Ce^{\large i\omega t}=C_1\cos\omega t+C_2\sin\omega t\]